OpenStudy (anonymous):

Determine how many grams of each of the following solutes would be needed to make 2.10x10^2 mL of a 0.100 M solution. (A) Cesium Iodide (CsI) (B) Sulfuric Acid (H2SO4) (C) Sodium Carbonate (Na2CO3) (D) Potassium Dichromate (K2Cr2O7) (E) Potassium permanganate (KMnO4) (D)

OpenStudy (anonymous):

every solution from A to E has the same mols n=MV=0.1*2.1*10^2*10^(-3)=0.021 mol

OpenStudy (anonymous):

but how many grams of each of the solutes would be needed u just need to find equivalent grams for 0.021 mol for example (B) Sulfuric Acid (H2SO4) grams=moles*Molecular weight=0.021*98=2.058 grams

OpenStudy (anonymous):

Thank you so much for the help!! I have alot more questions if you are on and have time could please help!